The number of 4 letter words can be formed from the letters of the word ‘INFINITY’ so that 2 are same 2 are different is:
 

Grade:12

1 Answers

subbu
5 years ago

Answer:

The Reqd. No. of Permutations=3360.

Explanation:

Suppose that, out of n things, r1 are of first type, r2 are of

second type, r3 are of third type,..., where r1+r2+r3+...=n.

Then, no. of possible distinct permutations is given by

n!(r1!)(r2!)(r3!)...

In our Example, there are total 8 letters in the word INFINITY ,

out of which, 3 letters are of one type (i.e., the letter I ), 2

are of second type (i.e., the letter N ) and the remaining 3 are

(i.e., the letters F,T and Y) are each of 1 type.

Thus, n=8,r1=3,r2=2,r3=r4=r5=1.

The Reqd. No. of Permutations=8!(3!)(2!)(1!)(1!)(1!)

=8×7×6×5×42!=3360.

i think this is the answer